OpenStudy (anonymous):

Find the direct variation equation of the graph through the points (0, 0) and (1, -2). Write in y=kx form. A.y = 2x B.y = -2x C.y = 1/2 x D.y = -1/2 x Find the direct variation equation which passes through (0, 0) and (4, 1). A.y = 4x B.y = -4x C.y = 1/4 x D.y = -1/4 x

jimthompson5910 (jim_thompson5910):

y = kx -2 = k(1) -2 = k k = -2 So the equation is y = -2x ------------------------------------ y = kx 1 = k(4) 1/4 = k k = 1/4 So the equation is y = 1/4 x

OpenStudy (campbell_st):

find the gradient m = rise/run. rise is -2, run is 1..... therefore m = -2 y-intercept is 0.... therefore the equation is y = -2x

OpenStudy (anonymous):